OpenStudy (anonymous):

Find the equation of the line that passes through (5,0) and has a slope of -2/3

OpenStudy (anonymous):

You want to start with point-slope form. (y-y1)=m(x-x1)

OpenStudy (anonymous):

use slope intercept form y=mx+c where m=-2/3 y=0 x=5 put in the y=mx+c to get c.

OpenStudy (anonymous):

Since (5,0) is a y-intercept, you can put it in the form y = mx + b. m = slope b = y=intercept So y = -2/3x + 5